# A model-agnostic gravitational-wave background characterization algorithm

### v1

@@ V.2 Foreground event subtraction
In this section we consider recovery of the unresolved background of BNSs after resolved signals have been subtracted. Subtraction is performed by removing contributions from resolvable CBCs, astrophysical backgrounds, and cosmological energy densities from $\Omega_{\textrm{GW}}$, with the remaining portion of $\Omega_{\textrm{GW}}$ constituting the signal we characterize [33, 99, 100, 101, 32, 102, 103, 104]. With next-generation detectors, it is expected that a fraction of BNSs and all but the furthest BBHs are resolvable, so the remaining unresolved signals will produce a GWB [105, 93, 106, 107, 108] and will be largely unresolvable due to confusion noise [109]. Therefore, in stochastic analyses with third-generation detector data, we are likely to fit the residual BNS signal after subtracting louder foreground events.
@@
For the reasons discussed above, we choose to inject the $\Omega_{\rm BNS,residual}$ curve from Fig. 1e of [33], which is the residual signal for BNSs for an SNR threshold of 8.0 in the HLV 3G three-detector network after subtracting foreground and resolvable noise sources. This signal is interesting because it has a high frequency turnover similar to the CBC injection from the previous section. However, $\Omega_{\rm BNS,residual}$ also has a low-frequency feature due to the limited detector sensitivity at low frequencies.
@@
Figure 4: Recovered posteriors of the unresolved BNS signal, $\Omega_{\rm BNS,residual}$ after subtracting resolved signals with SNR$>8$ using CE noise. The injected signal we use is from [33]. The left plots shows the recovered posteriors for $N_{\rm max}=30$ and the top right plots shows the recovered posteriors for $N_{\rm max}=2$. Each color envelope represents 95% confidence for the recovered posteriors for various observation times for the simulated CE noise (legend). The bottom plots provide the residuals, $\Omega_{\rm BNS,residual;res}$, for the 95% confidence envelopes subtracted from the injected signal, $\Omega_{\rm BNS,residual}$, from the top row. Limiting $N_{\rm max}=2$ reveals PL posteriors that track the injection at later frequencies, while allowing $N_{\rm max}=30$ detects a low-frequency subtraction feature. Mostly positive residuals $\Omega_{\rm BNS,residual;res}$ indicate that fitting a non-PL signal with a PL biases the fit, causing posterior draws miss sections of the injection and consequentially underfit the PL portion.
@@
The population of BNSs used in [33] is generated using a Monte Carlo procedure and fiducial parameter distributions [110, 92, 111, 112]. The full $\Omega_{\rm BNS}$ signal is the superposition of 1,438,835 BNSs for one year of observation using an SFR-informed redshift distribution with local BNS merger rate $R(z=0)=920\,{\rm Gpc}^{-3}{\rm yr}^{-1}$ [113], intrinsic source frame masses drawn from $\mathcal{N}(1.33\,{\rm M}_{\odot},0.09\,{\rm M}_{\odot})$, and BNS sky locations and angular parameters drawn from uniform distributions. We refer the reader to [33] for more details on the BNS population used to generate the full $\Omega_{\rm BNS}$ signal that we obtain $\Omega_{\rm BNS,residual}$. Although more detailed Monte Carlo analyses that quantify the strength of the unresolved background using more recent population models and rates have been performed since those in [33] (see, e.g. [46]), we proceed with $\Omega_{\rm BNS,residual}$ for its notable low-frequency feature relevant to this study.
[…]
@@
We now choose $N_{\rm max}=2$, probing the effects of fitting a non-PL signal with a PL signal using our model. To evaluate the effects of applying a PL to $\Omega_{\rm BNS,residual}$, we implement linear interpolation in the posteriors, which are shown for the same observation times as for $N_{\rm max}=30$ (top right) along with the residuals between the injected signal and the posteriors, $\Omega_{\rm BNS,residual;res}$ (bottom right). The PLs recovered in the top right panel of Fig. 4 are consistently shallower than the PL portion of the injection. The residuals between the signal and the posteriors confirm this, as the 95% confidence envelopes on the residuals are mostly positive. Therefore, by trying to fit an incorrect or overly constrained model to an injection using this transdimensional spline interpolation model, we will obtain posteriors with a bias towards the ignored features, leading to a recovery that is not truly accurate anywhere. In fact, fixing a PL model as sensitivity increases may risk recovery bias with any potential recovery algorithm, not just with this spline interpolation model. So, implementing this spline interpolation model with its full flexibility is the best way to consistently and correctly recover an injected signal.
V.3 First order phase transition-informed BPL
@@
Figure 5: Results for injecting and recovering a FOPT-informed smoothed BPL with a CE sensitivity curve. The injected signal is described in Eq. (14) from [33]. We show the 95% confidence envelopes for the recovered posteriors for observation times ranging from 1 week to 2 years on the left. The injected signal is overlaid (red dashed). On the right, we show a heatmap of the number of knots in each posterior (color bar) as a function of observation time. As observation time increases, we can better constrain the recovered posteriors.
@@
Figure 6: Violin plots showing the recovered turnover frequency as a function of observation time for injected and recovered FOPT-informed smoothed BPL with CE sensitivity. These turnover frequencies are computed from the same posteriors whose results are summarized in Fig. 5. For each posterior and observing time, we compute the frequency where $\Omega_{\rm GW}$ obtains its maximum, deeming that the turnover frequency. For observation times at a month and shorter, the violin plots are multimodal with peaks around $36$ Hz and $13$ Hz, which correspond to posteriors that do not turn over. As observation time increases, the posterior turnover frequencies converge at the injected frequency (black dashed line).
@@
We next inject a smoothed BPL, a common signal shape that arises when predicting the shape of a primordial GWB spectrum [114, 16, 36, 115]. Specifically, a BPL is a reasonable fit to the predicted spectrum from first-order phase transitions (FOPT) in the early Universe. It is also a crude model for the stellar-mass CBC background, which is a PL that turns over at high frequencies.
@@
FOPTs in the early universe involve the formation of lower-temperature bubbles of different phases of matter or fields that expand into their surrounding medium. FOPTs produce GWs through three main mechanisms: (1) bubble collisions that generate shock waves that produce GWs [18]; (2) sound waves in the plasma from bubble expansion, which are probably the dominant FOPT GW source [116]; and (3) plasma turbulence induced by bubble motion, especially during strongly first-order electroweak transitions [16].
@@
There are many analytical and numerical models for each FOPT effect that produces GWs. As in [36], we choose to model the frequency spectrum as a smoothed BPL:
@@
As in [36], we use parameters $\alpha_{1}=3,\alpha_{2}=-4,\Delta=2,f_{*}=30,$ and $\Omega_{*}=1.8\times 10^{-10}$.
@@
We show the recovery of the BPL turnover frequency, $f_{\rm turn}$, in Fig. 6. Each violin is a histogram of the $f_{\rm turn}$ for 50,000 posterior samples for each $T_{\rm obs}$. The ends of the violins are the extrema for each recovery histogram. We find that we can recover $f_{\rm turn}$ that peak near the injected turnover (black dashed line) for each observation time. For one month or less of observation time, the histograms have a peak beyond $35$ Hz, as well as a peak near the injected $f_{\rm turn}$. This means that some of the posteriors are PLs or diverge for shorter observation times. As the observing time increases, the posteriors peak near the injected $f_{\rm turn}$ and the extrema narrow around the injected value.
Similarly to Section V.1, we can examine the distribution of the number of knots used in the recovered posteriors. At only 1 week of observing time, posteriors have anywhere from 1 to 4 knots, but as we increase the observing time, we obtain posteriors with mostly 3 knots. This shift occurs after 30 days of observation time, which matches the behavior of the posterior envelope in the left panel of Fig. 5. We can conclude that the fitter opts for the minimum required number of knots to produce an accurate posterior for the injected signal.
@@ VI Constraining the merger rates of CBCs
With the increasing number of CBC detections, we can better determine the properties of their underlying populations, and ultimately study the populations of their massive star progenitors [117, 118]. Past work has illustrated a joint hierarchical analysis involving resolved and unresolved signals to place constraints on the merger rate of black holes as a function of redshift [40]. Recently, new methods have been proposed to constrain the GWB from black hole binaries in dense stellar clusters [13], and to constrain the metallicity dependence of black hole formation and the distribution of time delays between black hole formation and merger [52].
@@
Figure 7: A flowchart for the hierarchical implementation of the spline interpolation model in Sec. VI. Blue (orange) boxes indicate actions that occur in the $\mathcal{R}(z)$ ($\Omega_{\rm GW}(f)$) space and the green box is where the sampler returns the returns of the RJMCMC sampling of the spline interpolation model. Boxes in the gray region are the steps that occur within the hierarchical sampling and model implementation. By moving between the $\mathcal{R}(z)$ and $\Omega_{\rm GW}(f)$ spaces in the sampler, we can perform hierarchical fitting using our transdimentional spline interpolation model.
@@
Figure 8: Results for injecting an $\Omega_{\rm GW}$ and recovering both the $\Omega_{\rm GW}$ and its hyperparameter $\mathcal{R}(z)$ from Eq. (11). On the left, we show the recovery envelopes for an injected $\mathcal{R}(z)$ (black) with 68%, 95%, and 99.7% confidence envelopes (blue) corresponding to 1, 2, and 3 standard deviations. The right plot shows the 95% confidence recovery envelope (blue) for the corresponding $\Omega_{\rm GW}$ (black) with CE sensitivity (red). Although the posteriors have deviations from the smooth injected $\mathcal{R}(z)$, we still recover the correct turnover shape of the injected Madau-Dickinson-informed $\mathcal{R}(z)$ and the correct $\Omega_{\rm GW}$ using our transdimensional spline interpolation algorithm.
@@
Here, we employ this transdimensional algorithm to directly infer a BPL merger rate model, as in Eq. (13). This approach demonstrates that hierarchical analysis can be conducted effectively on a detected background signal with next-generation detectors, thereby revealing further insights that can be extracted from an observed GWB. For this merger rate mode, we utilize parameters: $R_{0}=25\,{\rm Gpc}^{-3}{\rm yr}^{-1}$, $\alpha=2.7$, $\beta=3.4$, and $z_{p}=2.4$ [40]. In this injection, we implement the same CE sensitivity curve used in Section V with 12 hours of observing time.
@@
We point to Fig. 7 for the workflow of how we inject an $\Omega_{\rm GW}(f)$ signal and utilize our spline interpolation model to recover the true $\mathcal{R}(z)$ associated with the $\Omega_{\rm GW}(f)$ injection. First, similar to the examples in Sec. V, we place the interpolating knots in redshift space and implement a few constraints. The knot $z$-value priors are placed linearly in the redshift space with a maximum of 30 knots that range in placement from $z=0$ to $z=10$, and their amplitudes are log-uniformly distributed between 0.1 and $10^{3}$. We do not allow knots to be placed at $z>10$ because we assume that no mergers occur earlier than $z=10$ [40]. Second, we do not fix the number of knots because we want to test the full flexibility of the model and its ability to recover signals with an efficient number of knots. However, we fix the amplitude of the first knot at $z=0$ because $\mathcal{R}(z=0)$ is astrophysically well constrained. The $\mathcal{R}(z)$ that we inject is a perfect BPL, and we would like to test whether the model can recognize the BPL and implement an appropriate number of knots.
@@
Because we apply the spline interpolation model in the redshift space and not in the frequency space, the recovered $\Omega_{\rm GW}(f)$ is determined differently than in the previous examples (Sec. V). In previous injection and recovery examples, we place the knots in the frequency space, which extracts the $\Omega_{\rm GW}(f)$ at each frequency independently from the adjacent frequencies. For the example in this section, we place the knots in the redshift space, which do not directly map to knots in the frequency space. For this reason, changes to the knots, and subsequently the interpolation model, in the redshift space induce changes to $\Omega_{\rm GW}(f)$ at a number of frequencies. This effect is not limited to our choice to fit $\mathcal{R}(z)$ with our spline interpolation model – we expect to see an equivalent effect if we were to instead fit the mass probability distribution (Eq. (V.1)). The successful hierarchical analysis of this example demonstrates the versatility of our model in fitting an $\Omega_{\rm GW}(f)$ signal across different sensitivity regimes and extracting crucial astrophysical information from signals.

### v2

@@ V.2 Foreground event subtraction
In this section, we consider the situation where resolvable foreground CBC events have been subtracted, and the unresolved GW background after this subtraction is the dominant contribution to the measured spectrum. Subtraction is typically performed by removing contributions from resolvable CBCs [Sachdev:2020bkk, Regimbau:2017, Harms_2008, Crowder_2005, Cutler:2005qq, Zhong:2022ylh, Zhong:2024dss, Zhong:2025qno], and we consider the scenario in  [Sachdev:2020bkk] where there is also some contribution to the signal from subtraction error. With next-generation detectors, it is expected that a fraction of BNSs and all but the furthest BBHs are resolvable, so the expected dominant contribution to the unresolved background, will likely be the residual BNS signal.
@@
For the reasons discussed above, we choose to inject the $\Omega_{\rm BNS,residual}$ curve from Fig. 1(e) of [Sachdev:2020bkk]. This is the combined signal from the unresolved BNS background and foreground subtraction error in the HLV 3G three-detector network. The curve corresponds to subtraction of foreground events that have a network SNR greater than 8. This signal is interesting because it has a high-frequency turnover similar to the CBC injection from the previous section, in addition to a low-frequency feature due to the limited detector sensitivity at those frequencies.
@@
Figure 5: Recovered posteriors of the unresolved BNS signal $\Omega_{\rm BNS,residual}$ after subtracting resolved signals with SNR>8>8 using CE noise. The injected signal we use is from [Sachdev:2020bkk]. The left plots shows the recovered posteriors for $N_{\rm max}=30$ and the top right plots shows the recovered posteriors for $N_{\rm max}=2$. Each color envelope represents 95% confidence for the recovered posteriors for various observation times for the simulated CE noise (legend). The bottom plots provide the residuals $\Omega_{\rm BNS,residual;res}$ for the 95% confidence envelopes subtracted from the injected signal, $\Omega_{\rm BNS,residual}$, from the top row. Limiting $N_{\rm max}=2$ reveals PL posteriors that track the injection at later frequencies, while allowing $N_{\rm max}=30$ detects a low-frequency subtraction feature. Mostly positive residuals $\Omega_{\rm BNS,residual;res}$ indicate that fitting a non-PL signal with a PL biases the fit, causing posterior draws miss sections of the injection and consequentially underfit the PL portion.
@@
The population of BNSs used in [Sachdev:2020bkk] is generated using a Monte Carlo procedure and fiducial parameter distributions [Regimbau:2016ike, Regimbau:2011rp, Meacher_2015, abbottGW170817ImplicationsStochastic2018]. The full $\Omega_{\rm BNS}$ signal is the superposition of 1,438,835 BNSs for one year of observation using an SFR-informed redshift distribution with local BNS merger rate $R(z=0)=920\,{\rm Gpc}^{-3}{\rm yr}^{-1}$ [Abbott_2019], intrinsic source frame masses drawn from $\mathcal{N}(1.33\,M_{\odot},0.09\,M_{\odot})$, and BNS sky locations and angular parameters drawn from uniform distributions. We refer the reader to [Sachdev:2020bkk] for more details on the BNS population used to generate the full $\Omega_{\rm BNS}$ signal that we obtain $\Omega_{\rm BNS,residual}$. Although more detailed Monte Carlo analyses that quantify the strength of the unresolved background using more recent population models and rates have been performed since those in [Sachdev:2020bkk] (see, e.g., [Bellie:2023jlq]), we proceed with $\Omega_{\rm BNS,residual}$ for its notable low-frequency feature relevant to this study.
We inject the signal $\Omega_{\rm BNS,residual}$ with a CE sensitivity curve, assuming $T_{\rm obs}=1,\,5,\,20\,\textrm{yr}$, and present the results of these recoveries in Fig. 5. To start, we allow $N=30$, which provides the interpolation model with the most flexibility. We maintain linear interpolation in log space between the recovered knots. The results of the 95% confidence envelope for a range of $T_{\rm obs}$ (legend) are shown in the left panel of Fig. 5. For each of the observation times, the posteriors (top left) recover the PL behavior from $\sim 15$-$60$ Hz and detect the upturn at $\sim 10$ Hz. As the observation time increases, the lower bounds of the 95% confidence envelopes align more closely with the injection, while the upper bounds effectively detect the injected signal for each observation time. In general, increasing the maximum number of knots allows the posteriors to recover more features with the trade-off of overfitting regimes of less sensitivity. For example, at high frequencies, where we have less sensitivity, the 95% confidence envelopes are wider.
@@
We now choose $N_{\rm max}=2$, probing the effects of fitting a non-PL signal with a PL signal using our model. To evaluate the effects of applying a PL to $\Omega_{\rm BNS,residual}$, we implement linear interpolation in the posteriors, which are shown for the same observation times as for $N_{\rm max}=30$ (top right) along with the residuals between the injected signal and the posteriors, $\Omega_{\rm BNS,residual;res}$ (bottom right). The PLs recovered in the top right panel of Fig. 5 are consistently shallower than the PL portion of the injection. The residuals between the signal and the posteriors confirm this, as the 95% confidence envelopes on the residuals are mostly positive. Therefore, by trying to fit an incorrect or overly constrained model to an injection using this transdimensional spline interpolation model, we will obtain posteriors with a bias toward the ignored features, leading to a recovery that is not truly accurate anywhere. In fact, fixing a PL model as sensitivity increases may risk recovery bias with any potential recovery algorithm, not just with this spline interpolation model. So, implementing this spline interpolation model with its full flexibility is the best way to consistently and correctly recover an injected signal.
V.3 First-order phase transition-informed BPL
@@
Figure 6: Results for injecting and recovering a FOPT-informed smoothed BPL with a CE sensitivity curve. The injected signal is described in Eq. (14) from [Sachdev:2020bkk]. We show the 95% confidence envelopes for the recovered posteriors for observation times ranging from 1 week to 2 years on the top left. The injected signal is overlaid (black solid line). The bottom left panel shows the residuals, $\Omega_{\rm GW,res}$, which are obtained by subtracting the 9595% confidence intervals in the top left panel from the injected signal. On the right, we show a heat map of the number of knots in each posterior (color bar) as a function of observation time. As observation time increases, we can better constrain the recovered posteriors.
@@
Figure 7: Violin plots showing the recovered turnover frequency as a function of observation time for injected and recovered FOPT-informed smoothed BPL with CE sensitivity. These turnover frequencies are computed from the same posteriors whose results are summarized in Fig. 6. For each posterior and observing time, we compute the frequency where $\Omega_{\rm GW}$ obtains its maximum, deeming that the turnover frequency. For observation times at a month and shorter, the violin plots are multimodal with peaks around $36$ and $13$ Hz, which correspond to posteriors that do not turn over. As observation time increases, the posterior turnover frequencies converge at the injected frequency (black dashed line).
@@
We next inject a smoothed BPL, a common signal shape that arises when predicting the shape of a primordial GWB spectrum [Caprini:2007xq, Huber:2008hg, Martinovic:2020hru, Romero:2021kby]. Specifically, a BPL is a reasonable fit to the predicted spectrum from first-order phase transitions (FOPT) in the early Universe. It is also a crude model for the stellar-mass CBC background, which is a PL that turns over at high frequencies.
@@
FOPTs in the early universe involve the formation of lower-temperature bubbles of different phases of matter or fields that expand into their surrounding medium. FOPTs produce GWs through three main mechanisms: (i) bubble collisions that generate shock waves that produce GWs [Hindmarsh:2013xza]; (ii) sound waves in the plasma from bubble expansion, which are probably the dominant FOPT GW source [Weir_2018]; and (iii) plasma turbulence induced by bubble motion, especially during strongly first-order electroweak transitions [Huber:2008hg].
@@
There are many analytical and numerical models for each FOPT effect that produces GWs. As in [Martinovic:2020hru], we choose to model the frequency spectrum as a smoothed BPL:
@@
As in [Martinovic:2020hru], we use parameters $\alpha_{1}=3,\,\alpha_{2}=-4,\,\Delta=2,\,f_{*}=30,$ and $\Omega_{*}=1.8\times 10^{-10}$.
@@
We show the recovery of the BPL turnover frequency $f_{\rm turn}$ in Fig. 7. Each violin is a histogram of the $f_{\rm turn}$ for 50,000 posterior samples for each $T_{\rm obs}$. The ends of the violins are the extrema for each recovery histogram. We find that we can recover $f_{\rm turn}$ that peak near the injected turnover (black dashed line) for each observation time. For one month or less of observation time, the histograms have a peak beyond $35$ Hz, as well as a peak near the injected $f_{\rm turn}$. This means that some of the posteriors are PLs or diverge for shorter observation times. As the observing time increases, the posteriors peak near the injected $f_{\rm turn}$ and the extrema narrow around the injected value.
Similarly to Sec. V.1, we can examine the distribution of the number of knots used in the recovered posteriors. At only 1 week of observing time, posteriors have anywhere from one to four knots, but as we increase the observing time, we obtain posteriors with mostly three knots. This shift occurs after 30 days of observation time, which matches the behavior of the posterior envelope in the left panel of Fig. 6. We can conclude that the fitter opts for the minimum required number of knots to produce an accurate posterior for the injected signal.
@@ VI Constraining the merger rates of CBCs
With the increasing number of CBC detections, we can better determine the properties of their underlying populations, and ultimately study the populations of their massive star progenitors [Zevin:2017evb, Zevin:2020gbd]. Past work has illustrated a joint hierarchical analysis involving resolved and unresolved signals to place constraints on the merger rate of black holes as a function of redshift [Callister:2020arv]. Recently, new methods have been proposed to constrain the GWB from black hole binaries in dense stellar clusters [Kou:2024gvp], and to constrain the metallicity dependence of black hole formation and the distribution of time delays between black hole formation and merger [Turbang:2023tjk].
@@
Figure 8: A flowchart for the hierarchical implementation of the spline interpolation model in Sec. VI. Blue (orange) boxes indicate actions that occur in the $\mathcal{R}(z)$ $\Omega_{\rm GW}(f)$] space, and the green box is where the sampler returns the returns of the RJMCMC sampling of the spline interpolation model. Boxes in the gray region are the steps that occur within the hierarchical sampling and model implementation. By moving between the $\mathcal{R}(z)$ and $\Omega_{\rm GW}(f)$ spaces in the sampler, we can perform hierarchical fitting using our transdimensional spline interpolation model.
@@
Figure 9: Results for injecting an $\Omega_{\rm GW}$ and recovering both the $\Omega_{\rm GW}$ and its hyperparameter $\mathcal{R}(z)$ from Eq. (11). On the left, we show the recovery envelopes for an injected $\mathcal{R}(z)$ (black) with 68%, 95%, and 99.7% confidence envelopes (blue) corresponding to 1, 2, and 3 standard deviations, respectively. The right plot shows the 95% confidence recovery envelope (blue) for the corresponding $\Omega_{\rm GW}$ (black) with CE sensitivity (red). Although the posteriors have deviations from the smooth injected $\mathcal{R}(z)$, we still recover the correct turnover shape of the injected Madau-Dickinson-informed $\mathcal{R}(z)$ and the correct $\Omega_{\rm GW}$ using our transdimensional spline interpolation algorithm.
@@
Here, we employ this transdimensional algorithm to directly infer a BPL merger rate model, as in Eq. (13). This approach demonstrates that hierarchical analysis can be conducted effectively on a detected background signal with next-generation detectors, thereby revealing further insights that can be extracted from an observed GWB. For this merger rate mode, we utilize parameters $R_{0}=25\,{\rm Gpc}^{-3}{\rm yr}^{-1}$, $\alpha=2.7$, $\beta=3.4$, and $z_{p}=2.4$ [Callister:2020arv]. In this injection, we implement the same CE sensitivity curve used in Sec. V with 12 hours of observing time.
@@
We point to Fig. 8 for the work flow of how we inject an $\Omega_{\rm GW}(f)$ signal and utilize our spline interpolation model to recover the true $\mathcal{R}(z)$ associated with the $\Omega_{\rm GW}(f)$ injection. First, similar to the examples in Sec. V, we place the interpolating knots in redshift space and implement a few constraints. The knot $z$-value priors are placed linearly in the redshift space with a maximum of 30 knots that range in placement from $z=0$ to $z=10$, and their amplitudes are log-uniformly distributed between 0.1 and $10^{3}$. We do not allow knots to be placed at $z>10$ because we assume that no mergers occur earlier than $z=10$ [Callister:2020arv]. Second, we do not fix the number of knots because we want to test the full flexibility of the model and its ability to recover signals with an efficient number of knots. However, we fix the amplitude of the first knot at $z=0$ because $\mathcal{R}(z=0)$ is astrophysically well constrained. The $\mathcal{R}(z)$ that we inject is a perfect BPL, and we would like to test whether the model can recognize the BPL and implement an appropriate number of knots.
@@
Because we apply the spline interpolation model in the redshift space and not in the frequency space, the recovered $\Omega_{\rm GW}(f)$ is determined differently than in the previous examples (Sec. V). In previous injection and recovery examples, we place the knots in the frequency space, which extracts the $\Omega_{\rm GW}(f)$ at each frequency independently from the adjacent frequencies. For the example in this section, we place the knots in the redshift space, which do not directly map to knots in the frequency space. For this reason, changes to the knots, and subsequently the interpolation model, in the redshift space induce changes to $\Omega_{\rm GW}(f)$ at a number of frequencies. This effect is not limited to our choice to fit $\mathcal{R}(z)$ with our spline interpolation model–we expect to see an equivalent effect if we were to instead fit the mass probability distribution [Eq. (V.1)]. The successful hierarchical analysis of this example demonstrates the versatility of our model in fitting an $\Omega_{\rm GW}(f)$ signal across different sensitivity regimes and extracting crucial astrophysical information from signals.